The local theater has three types of seats for Broadway plays: main floor, balcony and mezzanine. Main floor tickets are $58, balcony tickets are $44, and mezzanine tickets are $37. One particular night, sales totaled $82,812. There were 79 more main floor tickets sold than balcony and mezzanine tickets combined. The number of balcony tickets sold is 333 more than 4 times the number of mezzanine tickets sold. How many of each type of ticket were sold?
Answer by stanbon(75887)

Main floor tickets are $58, balcony tickets are $44, and mezzanine tickets are $37. One particular night, sales totaled $82,812.
There were 79 more main floor tickets sold than balcony and mezzanine tickets combined.
The number of balcony tickets sold is 333 more than 4 times the number of mezzanine tickets sold.
How many of each type of ticket were sold?
----------
58m +44b + 37z = 82,812
m = b + z + 79
b = 4z + 333
========================
Rearrange:
58m +44b + 37z = 82,812
m -b - z = 79
0 +b - 4z = 333
--------------------------
m = 852
b = 685
z = 88
==============
Cheers,
Stan H.